We consider the non-homogeneous problem y = 30(18x – 2x4) First we consider the homogeneous problem y = 0 : 1) the auxilia
Answer #1

4 = 30 (184-2X) Y=5408–60x4 0 ① and bot(= r2=0 Answer Answer r=0,0 © 2 теtу rе o - Главен 3 faundamental solution set is 1

Arswen 9 Hence = -2X6 + 90 x3 Y=G+6x-ax®+goxs Put 4 (0) =-), 7 (0) = 1 >-1=( 1=Cg Herce (3) Y=-148-2x + 90 y3 answer